OpenStudy (marcelie):

help please solve the differential equation dy/dx = x^2y-y+x^2-1

OpenStudy (marcelie):

@mathmale

OpenStudy (marcelie):

@zepdrix

OpenStudy (marcelie):

i was thinking of factoring iit \[\frac{ dy }{ dx }= y(x^2-1)+(x^2-1)\]

zepdrix (zepdrix):

Ooo nice :) I didn't see that.

OpenStudy (marcelie):

\[dy/dx = (y+1)+(x^2-1)\]

zepdrix (zepdrix):

Woops!

zepdrix (zepdrix):

\[\large\rm \left[y\color{orangered}{(x^2-1)}+1\color{orangered}{(x^2-1)}\right]\quad=\quad \color{orangered}{(x^2-1)}[y+1]\]

zepdrix (zepdrix):

No plus between the factors, ya?

OpenStudy (marcelie):

oh crapt...yeh

OpenStudy (marcelie):

so then hmm

zepdrix (zepdrix):

Now sure? Separate variables. Get the y's on the left, leave the x's on the right.

OpenStudy (marcelie):

\[|(y+1)|dy=(x^2-1)dx\]

zepdrix (zepdrix):

Hmm I'm not sure what you did... How did you move (y+1) to the left side..?

zepdrix (zepdrix):

It was `multiplying` the x stuff, so to move it to the left you must _________.

OpenStudy (marcelie):

oh i did a mistake arghh okay so the 1 _____ dy = (x^2-1) dx (y+1)

zepdrix (zepdrix):

Good

OpenStudy (marcelie):

so then hmmm integerate

OpenStudy (marcelie):

ln | y+1| = 1/3 x^3-1x

OpenStudy (marcelie):

okay so then we put e